The sum of two numbers is 48. The difference between three times the smaller number and the larger number is 16. Find the numbers.
Answer by checkley75(3666)

X+Y=48 X=48-Y
X-3Y=16
(48-Y)-3Y=16
48-Y-3Y=16
-4Y=16-48
-4Y=-32
Y=-32/-4
Y=8 ANSWER
X+8=48
X=48-8
X=40 ANSWER
PROOF
40-3*8=16
40-24=16
16=16